Using polar coordinates, evaluate the integral ∫∫R sin(x^2+y^2)dA where R is the region 4≤x^2+y^2≤36.
1
Expert's answer
2020-11-09T13:36:25-0500


Let's draw this region:



Let's move on to polar coordinates

"x = r\cos \varphi ,\,\,y = r\sin \varphi ,\,\,0 < \varphi < 2\pi ,\,\,2 < r < 6"

Then

"\begin{array}{l}
\int\limits_R {\int {\sin \left( {{x^2} + {y^2}} \right)} } dA = \int\limits_0^{2\pi } {d\varphi } \int\limits_2^6 {r\sin {r^2}dr} = \frac{1}{2}\int\limits_0^{2\pi } {d\varphi } \int\limits_2^6 {\sin {r^2}d{r^2}} = - \frac{1}{2}\left. \varphi \right|_0^{2\pi }\left. {\cos {r^2}} \right|_2^6 = - \pi (\cos 36 - \cos 4) = \\
 = \pi (\cos 4 - \cos 36)
\end{array}"